OpenStudy (anonymous):

if (2^a)(2^b)=128, what is a+6 equal to? A.1 B.6 C. 7 d.12

OpenStudy (anonymous):

is there any other information!!!!!!

OpenStudy (anonymous):

or is the question correct!!!!!

OpenStudy (anonymous):

this can't be solved

OpenStudy (anonymous):

\(2^a+2^b=2^{a+b}=2^7 \implies a+b=7\). The givens are not enough to answer the question.

OpenStudy (anonymous):

I think you're asking about a+b. If so then \(a+b=7\), which is answer C.